A 65 year old employee retires from her employer with annual retirement payment received once per year of $60,000. Her remaining life expectancy is 15 years, and she has no surviving spouse

At a 4% discount rate, based on the risk of her employer paying her pension and the rate of similar risk investments, what is the lump-sum value of her pension at the time of her retirement?

The present value of annuity factor for n=15 and r=4% is 11.1184.
Answer #1

Retiree will get $60000 per year for next 15 years,

discount rate d = 4%

So, present value of this annuity using ordinary annuity formula is

PV = A*(1 - (1+d)^-t)/d = 60000*(1 - 1.04^-15)/0.04 = $667103.25

Or using present value of annuity factor for n=15 and r=4% is 11.1184 PV = 60000*11.1184 = $667104